### v1

@@ Abstract
Spoken language datasets are vital for advancing linguistic research, Natural Language Processing, and speech technology. However, resources dedicated to Italian, a linguistically rich and diverse Romance language, remain underexplored compared to major languages like English or Mandarin. This survey provides a comprehensive analysis of 66 spoken Italian datasets, highlighting their characteristics, methodologies, and applications. The datasets are categorized by speech type, source and context, and demographic and linguistic features, with a focus on their utility in fields such as Automatic Speech Recognition, emotion detection, and education. Challenges related to dataset scarcity, representativeness, and accessibility are discussed alongside recommendations for enhancing dataset creation and utilization. The full dataset inventory is publicly accessible via GitHub and archived on Zenodo, serving as a valuable resource for researchers and developers. By addressing current gaps and proposing future directions, this work aims to support the advancement of Italian speech technologies and linguistic research.
Keywords
@@
1 Introduction
1.1 Background and Motivation
Spoken language datasets play a pivotal role in linguistic research, natural language processing (NLP), and speech technology. They provide the foundational data necessary for tasks such as automatic speech recognition (ASR), speech synthesis, and sociolinguistic analysis. Despite significant advancements in the field, linguistic resources specific to Italian remain relatively limited, particularly in comparison to languages such as English and Mandarin. Italian, as a Romance language with rich phonological and dialectal diversity, offers unique challenges and opportunities for research. Its linguistic variability across regions, combined with socio-cultural nuances, makes it a compelling subject of study for developing robust speech models. The increasing demand for speech-based applications—ranging from virtual assistants to automated transcription systems—further underscores the importance of well-annotated and comprehensive Italian spoken language datasets. While several large-scale surveys and corpus compilations have been conducted for other languages, a focused survey detailing the variety, scope, and applications of spoken Italian datasets is notably absent, as to our current knowledge. This paper aims to address this gap by presenting a comprehensive overview of existing resources, their methodologies, and their contributions to the field.
1.2 Objectives of the Survey
@@
To ensure inclusivity, the survey includes both public and commercial datasets, focusing on those actively maintained as of November 2024 with sufficient available information. While dataset quality and availability were not independently verified, the survey synthesizes information provided by creators or documentation. By identifying gaps and limitations, it offers recommendations for creating future datasets, emphasizing collaboration and standardization to expand and improve spoken Italian corpora. While the paper highlights key datasets to illustrate the discussed categories and methodologies, the full collection comprises 66 datasets, encompassing a wide range of speech types, linguistic features, and applications. This complete dataset inventory is publicly accessible via GitHub and archived on Zenodo (DOI: 10.5281/zenodo.14246196), providing an useful and collaborative resource for researchers and developers.
1.3 Structure of the paper
@@
  • Section 6 outlines future directions and recommendations for addressing the identified gaps and improving dataset quality.
  • Section 7 concludes by summarizing the findings and emphasizing the importance of collaborative efforts for advancing spoken Italian datasets.
2 Categorization of Spoken Italian Datasets and Corpora
@@ 2.1.4 Read Speech
The MLS dataset [17, 18], for instance, features 279 hours of Italian read speech derived from public domain audiobooks. Another significant contribution comes from Mozilla’s Common Voice project [19], a crowdsourced effort that has amassed over 1,000 hours of Italian read speech contributed by volunteers from various demographics, ensuring a diverse representation.
2.2 By Source and Context
The source and context of speech datasets determine their applicability and potential use cases. Categorizing datasets by source sheds light on their relevance to specific domains and scenarios aligning the dataset with particular real world scenarios. Categorizing by context impacts the dataset applicability to use cases requiring specific conditions or settings (e.g. noisy environments, emotional speech, accents etc). From an acoustic point of view different sources (studio vs. field recordings) vary in signal-to-noise ratios, reverberation, and other acoustic properties while contexts such as background noise [20], speaker diversity, or overlapping speech may affect pre-processing and model training strategies e.g. [21].
2.2.1 Broadcast Media
@@
2.2.2 Telephone Speech
Telephone speech datasets are indispensable for developing telephony-based systems such as customer service automation, voice-based authentication, and call transcription tools [26]. The Italian Spontaneous Dialogue Telephony Speech Dataset - Nexdata [27] is a prominent resource, featuring 500 hours of conversational recordings made via mobile phones. These recordings include a balanced representation of regional accents and cover various familiar topics, simulating natural interactions in telephonic environments. Another notable dataset is the Italian Spontaneous Dialogue Dataset [3] by Defined.ai, which includes telephony-based conversations recorded under both noisy and silent conditions. This dataset’s emphasis on diverse environments makes it highly applicable to real-world telephony use cases. Additionally, the Aurora Project Database [28], collected over the GSM telephone network in vehicular environments, provides speech data influenced by background noise and motion, contributing to advancements in noise-robust telephony systems. Together, these resources support the development of telephony applications with robust performance across varied acoustic and contextual challenges.
2.2.3 Social Media and Online Platforms
@@ 2.2.4 Field Recordings
2.3 By Demographic and Linguistic Features
@@
Spoken Italian datasets vary significantly in their demographic and linguistic coverage. Datasets skewed toward a particular demographic (e.g., one gender, age group, or ethnicity) may lead to biased models, while including balanced demographics ensures models perform equitably across all groups, avoiding unfair outcomes. If instead the purpose is toward the study of a particular region or ethnic group, then it is worth noticing that the speech may be characterized by a different rhythm or intonation even within the same language, thus it is important to characterize these features to make the dataset useful in applications targeting a specific community. Also the lack of linguistic diversity in datasets leads to models that fail in real-world settings where varied accents or code-switching (mixing languages) are common. These distinctions highlight the sociolinguistic dimensions of speech data.
2.3.1 Regional Varieties and Dialects
@@
Datasets highlighting regional varieties and dialects are essential for preserving linguistic diversity and understanding the intricate patterns of variation within Italian. These datasets enable researchers to analyze phonological, syntactic, and lexical differences across regions and promote the development of language technologies that cater to regional linguistic characteristics. The VIVALDI corpus [31] is a cornerstone resource in this domain, offering extensive coverage of Italian dialects and minority languages such as Lombard, Sicilian, and Sardinian. By documenting speech from over 14 regions, this dataset supports studies in dialectology and sociolinguistics, as well as efforts in language preservation. The KIParla corpus [2, 1], while focused on standard Italian, includes contributions from speakers across various regions, reflecting a wide spectrum of regional accents and idiomatic expressions. This diversity makes it a valuable resource for analyzing how regional varieties influence spoken language in different contexts. Another noteworthy dataset is the LABLITA corpus [12, 13], which, in addition to its focus on spontaneous and natural speech, incorporates recordings from different Italian regions. This regional emphasis allows researchers to explore how standard Italian interacts with regional speech patterns in everyday communication. The DEMoS corpus [32, 33], while primarily centered on emotional speech, also includes instances of regional accents among its participants, providing an additional dimension of linguistic diversity. Similarly, the Common Voice project [19] features contributions from volunteers representing various regions of Italy, ensuring a broad geographic coverage that reflects the nation’s linguistic heterogeneity. Finally, the C-ORAL-ROM corpus [14], though multilingual, includes a substantial component of Italian speech from different regional contexts, enabling comparative studies of dialectal usage within the broader Romance language family.
2.3.2 Sociolinguistic Variation
@@
Sociolinguistic datasets capture variations in speech influenced by factors such as age, gender, socio-economic status, and educational background. These datasets are critical for understanding how social variables shape linguistic practices and for building inclusive language technologies. The KIParla corpus [2, 1] is a key resource for sociolinguistic research, featuring participants from diverse socio-economic, educational, and regional backgrounds. By including speakers of different ages and genders, this dataset allows for nuanced analyses of how social factors interact with linguistic features in Italian. The LABLITA corpus [12, 13] further enriches sociolinguistic research with its recordings of natural speech collected from various social contexts, such as face-to-face conversations and telephone calls. The dataset includes participants across a wide age range and from different social strata, offering a comprehensive view of spoken Italian as influenced by sociolinguistic diversity. The DEMoS corpus [32, 33], focused on emotional speech, also contributes to sociolinguistic studies by involving participants aged 20 to 64 from varied socio-economic backgrounds. This range supports the examination of how emotion and sociolinguistic variables interact in speech. Another significant dataset is the ITALIC corpus [34, 35], which features crowdsourced recordings from 70 speakers spanning different Italian regions, age groups, and genders. The diversity of contributors ensures broad coverage of sociolinguistic variation, making it suitable for research in intent classification and linguistic variation. Lastly, the Common Voice project [19] by Mozilla includes contributions from a wide demographic spectrum, with metadata detailing the age, gender, and accents of speakers. This sociolinguistic richness makes the dataset valuable for studying how demographic factors influence speech patterns in Italian.
2.3.3 Bilingual and Multilingual Corpora
@@
Bilingual and multilingual corpora enable cross-linguistic studies and support the development of multilingual speech technologies, such as machine translation, speech synthesis, and multilingual ASR systems. These datasets are particularly valuable for studying language contact, code-switching, and comparative linguistics. The Europarl-ST corpus [36, 37], derived from European Parliament debates, includes approximately 64.18 hours of Italian speech alongside translations into eight other European languages. This resource is instrumental for speech translation research, providing aligned data for tasks like machine learning and language modeling. The PortMedia French and Italian Corpus [38], focusing on tourist information and reservation dialogues, provides bilingual data that is highly relevant for domain-specific studies. With 604 Italian dialogues aligned with French data, this corpus supports applications in multilingual dialogue systems and domain-specific NLP. The MuST-C corpus [29], a multilingual speech-to-text dataset, provides high-quality audio-text pairs for automatic speech translation, including an Italian subset. This dataset is instrumental in advancing research on cross-linguistic speech technologies and translation systems. The M-AILABS Speech Dataset [39] also includes an Italian subset as part of its multilingual collection. Derived from public domain audiobooks, the dataset spans multiple languages, enabling comparative studies in read speech and providing training data for multilingual speech systems. The MULTEXT Prosodic Database [40] offers multilingual resources, including Italian, French, German, English, and Spanish. Its focus on prosodic annotations makes it particularly useful for studying intonation patterns across languages, supporting research in paralinguistics and language comparison. Lastly, the C-ORAL-ROM corpus [14] integrates Italian alongside other Romance languages, such as French, Spanish, and Portuguese. Its recordings of spontaneous speech in diverse settings allow for comparative studies of spoken language dynamics within the Romance language family.
3 Data Collection and Annotation Methodologies
3.1 Data Collection Techniques
@@
Effective data collection techniques are fundamental for creating robust spoken language datasets. These techniques involve selecting representative samples of speakers, designing recording protocols, and ensuring high-quality audio capture. The methods employed vary significantly depending on the dataset’s purpose, ranging from controlled studio settings to spontaneous field recordings. Proper planning in data collection ensures the resulting datasets meet the specific requirements for linguistic research, ASR training, or speech technology applications.
3.1.1 Recording Methods
@@ 3.1.2 Participant Selection
3.2 Annotation and Transcription Processes
@@ 3.2.2 Annotation Levels
3.2.3 Tools and Software Used
@@
3.3 Quality Control and Validation
@@
In Table 1, we summarize the characteristics of the documented datasets w.r.t. size, source and context, type of speech, demographic and linguistic features and transcription standard.
Table 1: Summary of Selected Datasets
@@
4 Applications of Spoken Italian Datasets
Spoken Italian datasets could serve as foundational resources across a range of disciplines, enabling advancements in both theoretical research and practical applications. Their diverse scope, encompassing regional varieties, demographic diversity, and speech types, allows for broad exploration in linguistics, speech technologies, education, and cultural studies. This section outlines key areas where these datasets may have a significant impact.
4.1 Linguistic Research
@@ 4.1.5 Discourse and Pragmatics
4.2 Natural Language Processing (NLP) and Speech Technologies
@@
4.2.1 Automatic Speech Recognition (ASR)
@@ 4.2.3 Machine Translation
4.2.4 Sentiment and Emotion Analysis
@@
4.2.5 Speech-Based Personalization

### v2

@@ Abstract
Spoken Italian datasets are curated collections of audio recordings featuring native Italian speech in various contexts (e.g., spontaneous dialogues, read text, telephone conversations), often accompanied by transcriptions or linguistic annotations. They serve as foundational resources for a wide range of applications, including Automatic Speech Recognition (ASR), Text-To-Speech (TTS) synthesis, emotion detection, and broader linguistic research. Despite Italian’s status as a richly diverse Romance language—marked by significant dialectal variation—publicly available large-scale corpora have remained comparatively underrepresented when contrasted with those of major world languages such as English or Mandarin.
In this survey, we present a comprehensive examination of 66 spoken Italian datasets, highlighting their key characteristics, data collection methodologies, and annotation frameworks. We categorize the datasets by speech type (e.g., conversational, monologic, spontaneous), by source (e.g., broadcast media, telephone calls, field recordings), and by demographic or linguistic attributes (including dialects and sociolinguistic features). Our analysis uncovers critical issues around dataset scarcity, demographic underrepresentation, and restricted accessibility, limiting broader research and development efforts. To address these gaps, we propose best practices and future directions—such as expanding demographic coverage, promoting open-access models, and standardizing annotation protocols—to enrich Italian speech data resources. The complete dataset inventory is publicly available on GitHub and archived on Zenodo, offering researchers and developers a valuable reference. By highlighting both the achievements and the shortcomings in existing resources, this work ultimately aims to foster collaboration and to spur further advancements in Italian speech technologies and linguistic research.
Index Terms:
@@
Index Terms:
[…]
=-15pt
I Introduction
I-A Background and Motivation
Spoken language datasets play a pivotal role in linguistic research, natural language processing (NLP) and speech technology. They provide the foundational data necessary for tasks such as automatic speech recognition (ASR), speech synthesis, and sociolinguistic analysis. Despite significant advancements in the field, linguistic resources specific to Italian remain relatively limited, particularly compared to languages such as English and Mandarin. Italian, as a Romance language with rich phonological and dialectal diversity, offers unique challenges and opportunities for research. Its linguistic variability across regions, combined with socio-cultural nuances, makes it a compelling subject of study for developing robust speech models. The increasing demand for speech-based applications, ranging from virtual assistants to automated transcription systems, further underscores the importance of well-annotated and comprehensive Italian spoken language datasets. Although several large-scale surveys and corpus compilations have been conducted for other languages, a focused survey detailing the variety, scope, and applications of spoken Italian datasets is notably absent, to the best of our knowledge. This paper aims to address this gap by presenting a comprehensive overview of existing resources, their methodologies, and their contributions to the field.
I-B Objectives of the Survey
@@
To ensure inclusivity, the survey includes both public and commercial datasets, focusing on those actively maintained as of November 2024 with sufficient available information. While dataset quality and availability were not independently verified, the survey synthesizes information provided by creators or documentation. By identifying gaps and limitations, it offers recommendations for creating future datasets, emphasizing collaboration and standardization to expand and improve spoken Italian corpora. While the paper highlights key datasets to illustrate the discussed categories and methodologies, the full collection comprises 66 datasets that encompass a wide range of speech types, linguistic features, and applications. This complete dataset inventory is publicly accessible via GitHub and archived on Zenodo (DOI: 10.5281/zenodo.14246196), providing an useful and collaborative resource for researchers and developers.
I-C Structure of the paper
@@
  • Section 6 outlines future directions and recommendations to address identified gaps and improve the quality of the data set.
  • Section 7 concludes by summarizing the findings and emphasizing the importance of collaborative efforts to advance spoken Italian datasets.
II Categorization of Spoken Italian Datasets and Corpora
@@ II-A4 Read Speech
The MLS dataset [17, 18], for instance, features 279 hours of Italian read speech derived from public domain audiobooks. Another significant contribution comes from Mozilla’s Common Voice project [19], a crowd-sourced effort that has collected over 1,000 hours of Italian read speech contributed by volunteers from various demographics, ensuring a diverse representation.
II-B By Source and Context
The source and context of speech datasets determine their applicability and potential use cases. Categorizing datasets by source sheds light on their relevance to specific domains and scenarios aligning the dataset with particular real world scenarios. Categorizing by context impacts the dataset applicability to use cases requiring specific conditions or settings (e.g. noisy environments, emotional speech, accents etc). From an acoustic point of view, different sources (studio vs. field recordings) vary in signal-to-noise ratios, reverberation, and other acoustic properties while contexts such as background noise [20], speaker diversity, or overlapping speech may affect pre-processing and model training strategies e.g. [21].
II-B1 Broadcast Media
@@
II-B2 Telephone Speech
Telephone speech datasets are indispensable to develop telephony-based systems such as customer service automation, voice-based authentication, and call transcription tools [26]. The Italian Spontaneous Dialogue Telephony Speech Dataset - Nexdata [27] is a prominent resource, featuring 500 hours of conversational recordings made via mobile phones. These recordings include a balanced representation of regional accents and cover various familiar topics, simulating natural interactions in telephonic environments. Another notable dataset is the Italian Spontaneous Dialogue Dataset [3] by Defined.ai, which includes telephony-based conversations recorded under both noisy and silent conditions. This dataset’s emphasis on diverse environments makes it highly applicable to real-world telephony use cases. Additionally, the Aurora Project Database [28], collected over the GSM telephone network in vehicular environments, provides speech data influenced by background noise and motion, contributing to advancements in noise-robust telephony systems. By leveraging these resources, developers can create telephony applications that maintain robust performance under diverse acoustic and contextual challenges.
II-B3 Social Media and Online Platforms
@@ II-B4 Field Recordings
II-C By Demographic and Linguistic Features
@@
Spoken Italian datasets can differ greatly in both demographic and linguistic coverage. When a dataset skews toward a specific demographic—such as a particular gender, age group, or ethnicity—it risks producing biased models. Conversely, ensuring balanced demographic representation helps models perform equitably across all groups, thereby preventing unfair outcomes. However, if the goal is to study a specific region or ethnic community, it is important to note that speech patterns can vary in rhythm and intonation even within the same language. Properly characterizing these features is essential for creating a dataset that serves the needs of targeted applications within that community. Also, the lack of linguistic diversity in datasets leads to models that fail in real-world settings where varied accents or code-switching (mixing languages) are common. These distinctions highlight the sociolinguistic dimensions of speech data.
II-C1 Regional Varieties and Dialects
@@
Datasets highlighting regional varieties and dialects are essential to preserve linguistic diversity and understand the intricate patterns of variation within Italian. These datasets enable researchers to analyze phonological, syntactic, and lexical differences across regions and promote the development of language technologies that cater to regional linguistic characteristics. The VIVALDI corpus [31] is a cornerstone resource in this domain, offering extensive coverage of Italian dialects and minority languages such as Lombard, Sicilian, and Sardinian. By documenting speech from over 14 regions, this dataset supports studies in dialectology and sociolinguistics, as well as efforts in language preservation. The KIParla corpus [2, 1], while focused on standard Italian, includes contributions from speakers across various regions, reflecting a wide spectrum of regional accents and idiomatic expressions. This diversity makes it a valuable resource for analyzing how regional varieties influence spoken language in different contexts. Another noteworthy dataset is the LABLITA corpus [12, 13], which, in addition to its focus on spontaneous and natural speech, incorporates recordings from different Italian regions. This regional emphasis allows researchers to explore how standard Italian interacts with regional speech patterns in everyday communication. The DEMoS corpus [32, 33], while primarily centered on emotional speech, also includes instances of regional accents among its participants, providing an additional dimension of linguistic diversity. Similarly, the Common Voice project [19] features contributions from volunteers representing various regions of Italy, ensuring a broad geographic coverage that reflects the nation’s linguistic heterogeneity. Finally, the C-ORAL-ROM corpus [14], though multilingual, includes a substantial component of Italian speech from different regional contexts, enabling comparative studies of dialectal usage within the broader Romance language family.
II-C2 Sociolinguistic Variation
@@
Sociolinguistic datasets capture variations in speech influenced by factors such as age, gender, socio-economic status, and educational background. These datasets are critical for understanding how social variables shape linguistic practices and for building inclusive language technologies. The KIParla corpus [2, 1] is a key resource for sociolinguistic research, featuring participants from diverse socio-economic, educational, and regional backgrounds. By including speakers of different ages and genders, this dataset allows for nuanced analyses of how social factors interact with linguistic features in Italian. The LABLITA corpus [12, 13] further enriches sociolinguistic research with its recordings of natural speech collected from various social contexts, such as face-to-face conversations and telephone calls. The dataset includes participants across a wide age range and from different social strata, offering a comprehensive view of spoken Italian as influenced by sociolinguistic diversity. The DEMoS corpus [32, 33], focused on emotional speech, also contributes to sociolinguistic studies by involving participants aged 20 to 64 from vaious socio-economic backgrounds. This range supports the examination of how emotion and sociolinguistic variables interact in speech. Another significant dataset is the ITALIC corpus [34, 35], which features crowd-sourced recordings from 70 speakers spanning different Italian regions, age groups, and genders. The diversity of contributors ensures broad coverage of sociolinguistic variation, making it suitable for research in intent classification and linguistic variation. Lastly, the Common Voice project [19] by Mozilla includes contributions from a wide demographic spectrum, with metadata detailing the age, gender, and accents of speakers. This sociolinguistic richness makes the dataset valuable for studying how demographic factors influence speech patterns in Italian.
II-C3 Bilingual and Multilingual Corpora
@@
Bilingual and multilingual corpora enable cross-linguistic studies and support the development of multilingual speech technologies, such as machine translation, speech synthesis, and multilingual ASR systems. These datasets are particularly valuable for studying language contact, code-switching, and comparative linguistics. The Europarl-ST corpus [36, 37], derived from European Parliament debates, includes approximately 64.18 hours of Italian speech alongside translations into eight other European languages. This resource is instrumental for speech translation research, providing aligned data for tasks as machine learning and language modeling. The PortMedia French and Italian Corpus [38], focusing on tourist information and reservation dialogues, provides bilingual data that is highly relevant for domain-specific studies. With 604 Italian dialogues aligned with French data, this corpus supports applications in multilingual dialogue systems and domain-specific NLP. The MuST-C corpus [29], a multilingual speech-to-text dataset, provides high-quality audio-text pairs for automatic speech translation, including an Italian subset. This dataset is instrumental in advancing research on cross-linguistic speech technologies and translation systems. The M-AILABS Speech Dataset [39] also includes an Italian subset as part of its multilingual collection. Derived from public domain audiobooks, the dataset spans multiple languages, enabling comparative studies in read speech and providing training data for multilingual speech systems. The MULTEXT Prosodic Database [40] offers multilingual resources, including Italian, French, German, English, and Spanish. Its focus on prosodic annotations makes it particularly useful for studying intonation patterns across languages, supporting research in paralinguistics and language comparison. Lastly, the C-ORAL-ROM corpus [14] integrates Italian alongside other Romance languages, such as French, Spanish, and Portuguese. Its recordings of spontaneous speech in diverse settings allow for comparative studies of spoken language dynamics within the Romance language family.
III Data Collection and Annotation Methodologies
III-A Data Collection Techniques
@@
Robust spoken language datasets depend on effective data collection strategies, which typically involve selecting representative speaker samples, designing appropriate recording protocols, and ensuring high-quality audio capture. The methods used can vary widely based on the dataset purpose, ranging from controlled studio environments to spontaneous field recordings. Proper planning in data collection ensures the resulting datasets meet the specific requirements for linguistic research, ASR training, or speech technology applications.
III-A1 Recording Methods
@@ III-A2 Participant Selection
III-B Annotation and Transcription Processes
@@ III-B2 Annotation Levels
III-B3 Tools and Software Used
@@
III-C Quality Control and Validation
@@
In Table I, we summarize the characteristics of the documented datasets w.r.t. size, source and context, type of speech, demographic and linguistic features and transcription standard.
TABLE I: Summary of Selected Datasets
@@
IV Applications of Spoken Italian Datasets
Spoken Italian datasets could serve as foundational resources across a range of disciplines, enabling advancements in both theoretical research and practical applications. Their diverse scope, encompassing regional varieties, demographic diversity, and speech types, allows for broad exploration in linguistics, speech technologies, education, and cultural studies. This section outlines key areas where these datasets may have a significant impact. In table II we summarize the relevance of 23 datasets w.r.t. the application fields outlined in this section.
IV-A Linguistic Research
@@ IV-A5 Discourse and Pragmatics
IV-B Natural Language Processing (NLP) and Speech Technologies
@@
IV-B1 Automatic Speech Recognition (ASR)
@@ IV-B3 Machine Translation
IV-B4 Sentiment and Emotion Analysis
@@
IV-B5 Speech-Based Personalization